### v1

# The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius

Abstract.
Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation by, e. g., improved data recording technologies and virtual reality devices. Immersive environments provide novel ways to present historical events or objects with high visual quality for both, the general public as well as researchers. Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics, utilising reconstructions of the mausoleum of the late Roman emperor Maxentius in Rome for argumentation. Based on our work, we advocate the value of combining historical and computer science expertise for virtual reconstructions and immersive environments to facilitate deeper understanding and interactive exploration of archaeological data.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JOCCHjournalvolume: xjournalnumber: xarticle: xpublicationmonth: 1ccs: Applied computing Architecture (buildings)ccs: Human-centered computing Visualization application domainsccs: Applied computing Computers in other domainsccs: Human-centered computing Computer supported cooperative workccs: Applied computing Archaeology
1. Introduction
Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualisation for work on cultural heritage (Reilly, 1990; Prescott, 1997). Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen immense development in its technological capabilities in terms of data acquisition and data representation, as well as several definitions of charters which employ meaningful standards for visualisations and virtual reconstructions (Lopez-Menchero and Grande, 2011; Denard, 2012). According to the Seville Charter (Lopez-Menchero and Grande, 2011), Virtual Archaeology concerns itself with finding comprehensive ways of visualising cultural heritage objects.
The visualisation of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e. g., models in virtual reality), since 2D visualisations often fall short of capturing the complexity and spatial relationships inherent in archaeological data. Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino (Milgram and Kishino, 1994). Most commonly, immersive technologies are understood to be either virtual reality (VR) as an entirely virtual environment, augmented reality (AR) as an overlay of additional information into the real world, or mixed reality (MR) as a blend between the virtual and real world. By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artefacts but also reconstructed objects, background information, or a contextually appropriate environment. This leads to a more immersive and interactive exploration of the past.
Rome, as a focal point of archaeological work in Europe, enjoys a high popularity for such work, leading to an increase in more extensive reconstruction visualisations. An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD (Frischer and Massey, 2022, pp. 23–25). However, since such reconstructions target a broader audience, their representations are often suggestive. The transparency of their underlying archaeological data origin tends to be selective, and their authenticity and reliability are therefore limited. This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction (Liverani, 2021, pp. 13–14).
However, to support scientific debates, digital reconstructions can be used to develop innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualise different chronological phases of archaeological sites, and reflect on historical monuments’ construction history and building process. A clear indication of different phases, information sources or certainty of validity can, among others, take the form of colour mappings or opacity of the respective components integrated into the visualisation (Bartz et al., 2016; Demetrescu, 2018; Pietroni and Ferdani, 2021).
To combine the potential of immersive environments and scientific debates, the relatively young field of Immersive Analytics (IA) offers plenty of advancements. IA uses engaging and embodied tools to support data understanding (Marriott et al., 2018), explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas. Therefore, IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality. As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it (Bienroth et al., 2022; Nowke et al., 2013). The investigation of which navigation or interaction methods are most appropriate for a specific context (Drogemuller et al., 2018), how effective VR is in teaching contexts (Sari et al., 2023), or the impact of different modalities like audio or haptics (Ruotolo et al., 2013; Venkatesan et al., 2023) are some of its core questions. The answers to those questions are as relevant for cultural heritage dissemination and research as they are for other applications. They support understanding of findings and to develop hypotheses, which in a reconstruction context means deciding on more meaningful interpretations of the original historical intent and use. In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, like annotating specific objects or positions or recording workflows. Creating immersive environments that allow people to explore sites, artefacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects, supports bridging the gap between academia and the public. Current scientific debates can be communicated by condensing the underlying information and providing a more relatable and captivating experience.
Since immersive experiences are broadly utilised in the context of dissemination, like museums or teaching, as well as for research (Boboc et al., 2022; Rodriguez-Garcia et al., 2024), and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies (Münster, 2019), we discuss perspectives of how the mentioned potentials of IA can be utilised. We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions, and utilise three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig. 1). Those reconstructions are based on ongoing scientific debates on what the mausoleum might have looked like in its prime, and we provide those in addition to software to inspect them in VR. Thus, our contribution is a perspective on where those two research directions can be merged to further discourse and show, by example, some of the possible concepts and techniques.
Maxentius A                          Maxentius B                        Maxentius C
Figure 1. This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work. The first model (Maxentius A) closely follows Rasch’s reconstruction proposal (Rasch, 1984). The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi (Johnson, 2009). The third model (Maxentius C) tries to merge the dome and the windows of Tor de’ Schiavi to the proportions of the Maxentius Mausoleum.
[…]
Renderings of all three digital reconstruction alternatives we propose next to each other with a focus on the different roofs, which pose the main difference between those models.
2. Background
Given the interdisciplinary nature of the topic, this section will introduce key vocabulary that is contextually relevant. In the work of Lopez-Menchero et al. (Lopez-Menchero and Grande, 2011), virtual reconstruction is defined as using a virtual model to recover a building or object made by humans visually. Here, we differentiate between digital reconstructions, virtual reconstructions, and immersive reconstructions to indicate the different contexts. Digital reconstruction is the process of digitally reconstructing an object through digital (computational) tools. Virtual reconstruction is the representation of a cultural heritage object in a virtual space to explore and interact with it. It is an application of digital reconstruction. Immersive reconstruction is the representation of a cultural heritage object in an immersive environment, that is, an environment along Milgram’s reality-virtuality continuum. It can be virtual, augmented, or mixed and can include additional techniques, such as data physicalisation, to provide an immersive experience.
One term relevant for reconstructions is anastylosis as part of virtual restoration, which describes the process of using in-situ findings as a reference, where in-situ means that they are found at their original cultural heritage site or, more specifically, directly “where they belong”. Those findings are then placed at their original position in the reconstruction and utilised to infer further missing components. This workflow is thoroughly described and used as part of the Extended Matrix by Demetrescu et al. (Demetrescu, 2018), who use colour coding to represent the reliability between different components of a reconstruction from still existing in-situ components which reside at their original position to purely hypothetical parts.
As we will discuss later in this work, virtual and immersive reconstructions often consist not only of the cultural heritage object itself but also of an artificial environment which, to some degree, resembles the environment where the object was formerly to be found. This concept adheres to the idea of situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making (Marriott et al., 2018, p. 185). For example, it might overlay an existing object with analysis information about it and is, thus, tightly linked to AR applications. Embodiment relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point and click interaction as it is typical for desktop PCs (Marriott et al., 2018, p. 3, p. 101). That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell. The psychological immersion, also called engagement (Marriott et al., 2018, p. 10), is another important concept, as there is evidence that task performance and recall are boosted through heightened engagement (Makowski et al., 2017). The last concept to mention is collaborative immersive analytics. In IA, many problems addressed require more than one person to solve (Marriott et al., 2018, p. 221), leading to multiple people working in the same virtual environment. This includes the collaboration of multiple people from different physical locations and even at various points in time.
3. Immersive Analytics for Virtual Archaeology
The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching and even research in archaeological contexts has grown (Boboc et al., 2022; Rodriguez-Garcia et al., 2024). As we focus on reconstructions in this work, we consider digital reconstruction models as well as the data recordings, which lead to those being the data we work with and seek to understand. Consequently, the general intersection with IA is clear: We use current technologies to present our data, and we want to optimise the experience to further a user’s ability to understand the reconstructions.
To better understand data, some level of engagement is desirable, as there is evidence for an increased recall of data (Mark et al., 2003). This can be achieved by using immersive VR environments, which are popular in different outreach-related domains (Frischer and Massey, 2022; Schofield et al., 2018; Kyrlitsias et al., 2020). Such environments have proven to create a heightened sense of presence, as already argued by Kyrlitsias et al. (Kyrlitsias et al., 2020), which matches findings from IA. For example, Ragan et al. argue that the combination of s3D and head-tracking helps spatial judgment accuracy (Ragan et al., 2012). The analysis they conducted in their study looked into the different effects of stereovision, head-tracking and bodily movement by tasking their participants with context-related tasks, in which the participant’s error rate was almost halved when using stereovision and head-tracking. Nowadays, such analysis can easily be done for cultural heritage dissemination environments, as many modern head-mounted display devices have integrated telemetry hardware like eye-tracking, and all of them have either internal or external head-tracking. Museums, for instance, could collect data to see which points of interest in their exhibit attract visitors. In general, an investigation could be conducted on how people perceive any given environment and whether the scene is overloaded or achieves the desired recall of the presented information.
AR is also already seeing applications in virtual archaeology  (Tscheu and Buhalis, 2016). Such environments create a heightened sense of presence, which matches findings from IA that situated analytics supports reasoning (ElSayed et al., 2015) by embedding data into the physical world (Willett et al., 2017). Several papers on situated analytics cover pipelines for creating corresponding environments (Kalkofen et al., 2007; Willett et al., 2017; Fleck et al., 2023), leading to the conclusion that these pipelines can be used for virtual reconstruction models as well. One widespread use for AR is mapping digital information onto a physical referent, an approach described by the conceptual model by Willett et al. (Willett et al., 2017). Given such a physical referent is a heritage site or a part of it, reconstructed components can be acquired from an existing data set like digital twins, digital reconstructions or even just raw point clouds, and the actual physical object can be supplemented by a virtual representation visible in the AR system. For example, a proximity-based magical lense visualisation (Lee et al., 2023) could be used, where supplement information for an object is only shown in an embedded fashion given a certain proximity of the user to the object. Other known or inferred information can be overlayed apart from reconstructed components. This could be, for example, the approximate age of an illustration or an estimated purpose of an object, which could be calculated in the background, or the “Logical World”, as Willett et al.  call it. Another information that could be provided that way is the different types of archaeological evidence as described by Demetrescu et al. (Demetrescu, 2018), from actual residuals of the degraded objects to inferences from typologically similar objects or lore. Those types of evidence can be mapped to different levels of certainty, from well-preserved to purely hypothetical, and should be highlighted appropriately in reconstruction visualisations (Demetrescu, 2018; Schäfer, 2018; Pietroni and Ferdani, 2021).
While visualisation is essential, current immersive environment devices offer a wide variety of possibilities to utilise different modalities for input and output. One of which is haptics, which supports sense-making for priorly unseen data (Taher et al., 2017). Encoding information like material density or even the age of different objects into the intensity of haptic feedback are possible channels for additional information. Furthermore, data physicalisation, for instance, in the form of 3D prints, is already being used to provide a deeper understanding of existing information (Klein et al., 2022), which gives users the possibility of actually touching their data. Such physicalisations could, in turn, be used as a way of interacting with reconstructions. For example, an AR device could recognise the region a user is pointing at to provide information about it. Another multi-modal channel for in- and output is sound, and there are various ways of using sound to map data (sonification) (Madhyastha and Reed, 1995). Having access to realistic sound models can already help to interpret which purpose a specific room might have served (McBride, 2014). Since ambient sound can reduce distraction and improve presence (Kern and Ellermeier, 2020), the identification and application of contextually appropriate ambient sounds which are based on a user position (in the digital world) or activities, as well as attention control through specifically designed noises can be used to improve dissemination quality or aid validity assessment.
Further intriguing aspects of IA environments are the possibilities that arise from development kits such as Unity (Unity Technologies, 2023) or Unreal Engine (Epic Games, 2024), which come with working physics engines and multi-user capabilities. Through physics engines, concepts like lighting can be realistically simulated to allow for the investigation of their influence on reconstructions. Through multi-user capabilities, multiple people who do not necessarily have to be at the same place (distributed collaboration) can look at reconstructions and interact with them, which augments the understanding of users (Hill, 1982; Mark et al., 2003). By integrating user interfaces, either simple VR versions of desktop-like interfaces or more natural ones (Wigdor, 2011), users can furthermore be given the power to manipulate objects or the environment as a whole directly. Definitions and a more generalised approach towards interaction systems can be found in IA-related work (Büschel et al., 2018). Such interaction systems can be utilised and combined to reasonable manipulation methods of a cultural heritage object, such as interactively changing the lighting conditions, adding annotations or filters for components, or reassembling pieces of the reconstructions in a playful manner (Houska et al., 2024). With a selection of those considerations, we will further discuss the example of the Maxentius Mausoleum and how these considerations can be applied to an environment for it.
4. Maxentius Mausoleum
Figure 2. The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
[…]
A photo we took of the Maxentius Mausoleum in 2024. Visible are the remains of the crypt from the outside as well as the farmhouse.
To provide insight into the significance of our archaeological example, we briefly discuss its historical and typological embedding and connection to the reign of Maxentius. The latter is relevant, as after nearly a century, during which the Roman emperors had been virtually absent from Rome and spent most of their time campaigning at the frontiers of the empire, the short rule of Maxentius (r. 306-312) gave Rome again the status of a permanent imperial residence. Maxentius styled himself as the conservator of Rome (conservator urbis suae) and added several large buildings such as the Maxentius Basilica and the temple of Venus and Rome, thermal baths and an extension of the imperial palace on the Palatine Hill to the architectural ensemble of the city (Leppin and Ziemssen, 2007, pp. 12-13, 51, 74-82, 87-95).
On the city’s outskirts, at the third milestone of the Via Appia leading southwards, Maxentius built a residential complex, which encompassed a villa suburbana with an attached well-preserved circus and a monumental mausoleum directed towards the Via Appia. The mausoleum, which forms the core of the complex, is surrounded by a slightly offset portico of 99.50 by 85 meters, of which substantial remains (as high as 11 meters) can still be seen (Rasch, 1984, p. 14) (see Fig. 2). However, the preserved remains of the mausoleum give the modern visitor only a limited idea of its former importance and structure. Just the foundations of the original building remain, overbuilt partly by a farmhouse of the 18th century (Rasch, 1984, pp. 5–6). Inside the building, an ambulatory crypt around an overarching pillar with niches inside is well-preserved. The layout of the building includes a round podium with a diameter of 33.25 meters. Adjacent is a front porch partly overbuilt by a farmhouse with a length of 17.00 meters and a width of 22.68 meters. Covered by the farmhouse, a vaulted cellar of the front porch with a diameter of 9.10 by 16.33 meters remains. The still-existing crypt has a diameter of 23.54 meters, a central column 9.70 meters wide, and seven niches (Johnson, 2009, pp. 86–90).
The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD (Johnson, 2009, pp. 92–93). Its few remaining traces provide an excellent case study for the application of virtual archaeology. Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon. From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modelled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey (Rasch, 1993, p. 84). However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) (Rasch, 1993). This monument is located, like the Maxentius Mausoleum, on the outskirts of Rome, close to an imperial villa suburbana. It follows the design of a podium round building featuring a slightly smaller, tetrastyle front porch with a domed vault without an oculus but four windows in the clerestory. These different perspectives, into which Maxentius’ Mausoleum can be classified, result in different conceivable models for its reconstruction, which will be discussed below.
5. Immersive Analytics for the Maxentius Mausoleum
Regarding the Maxentius Mausoleum, we chose Rasch’s work as a foundation for a first reconstruction (Maxentius A), which closely follows his suggested measurements. The second model (Maxentius B) assumes a closed dome following Johnson’s proposal (Johnson, 2009, pp. 91-92) and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi (Rasch, 1993, p. 28). The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi (Rasch, 1993, p. 68) and the windows with a larger diameter.
Figure 3. Cross-faded version of Maxentius A. Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
[…]
A transparent version of a digital reconstruction based on Rasch’s work is put on a photogrammetry-based digital twin of the current state of the mausoleum.
This model visualises, as suggested by Johnson, a version of the Maxentius Mausoleum modelled closely on Tor de’ Schiavi (Johnson, 2009, p. 92). Consequently, model B can be considered a merger of models A and C. Thus, we have three reconstruction models of the same object (see Fig. 1), which mainly differ in the dome’s design and allow us to discuss which of the hypotheses about the part of the mausoleum is most plausible. We created the reconstructions with the 3D modelling software Blender. The components of the models consist of combined simple geometric forms, which were manipulated through manual node shifting to obtain necessary details. For each sub-component, once constructed from primitives, the measurements for those components were linearly scaled and applied. Thanks to the precise measurements we get from Rasch’s work, the models provide correct proportions, which helps us to obtain a fitting impression of the mausoleum’s former glory. In Fig. 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
Figure 4. The software we provide together with our models. The panel, the miniature, and the actual visualisation can be seen.
[…]
An image from a user’s perspective, where the user sees an in-scale reconstruction of the Maxentius Mausoleum. What the user also sees is a miniature of the mausoleum reconstruction on a cube as well as a small panel with buttons with the names Maxentius A, B, and C, referring to the three available reconstructions.
We provide software (see Fig. 4) that presents the three hypothesised models of the Maxentius mausoleum. The software provides a Virtual Reality environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index. Inside the immersive environment, the user will start in front of the stairway of Maxentius A. The user can move around, using the continuous movement metaphor and can snap-turn or move around freely, depending on the physical space available around them. When starting the software, in front of the user is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a colouring akin to the extended matrix by Demetrescu et al. The buttons can be touched by the user with their hands. When the user touches one of the buttons, the presented model will switch, and the user will get haptic feedback.
To the right of the panel is a miniature of the model. Users can hold their virtual hands inside it and grab it with a simple grabbing motion. Upon grabbing, the user can control with their wrist the position of the sun and thus the light’s angle of incidence on the model. There is a small yellow sphere indicating the relative position of the light source, which imitates the sun so that users do not have to search for the real sun’s position. Once the user reaches for the miniature, a vibration of the controller indicates that they can now grab to rotate the light source. Once the grabbing motion is performed, a slightly more intense and more prolonged vibration pattern means that the wrist’s rotation now manipulates the light source’s angle. Furthermore, users can place small squares to leave small randomly coloured markers at points of interest behind. Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need of additional hardware. We provide the software via a Zenodo upload at (Kerle-Malcharek, 2025).
6. Discussion
Our exploration of IA applications in Virtual Archaeology argues for a significant potential for enhancing data understanding and analysis in archaeological reconstructions. While there exist works which advocate for the usefulness of immersive environments in Virtual Archaeology (Kyrlitsias et al., 2020), many works focus on dissemination or reconstruction. Therefore, this work yields significance for the field by presenting general considerations of how research can benefit from immersive technologies.
For our considerations, we utilise the Maxentius Mausoleum in Rome as a case study in the form of three digital reconstructions. The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea, which we also present, and are based on archaeological evidence and scholarly interpretations. We did this since we wanted to use reconstructions based on scientific evidence and valid assumptions to provide a sound foundation for our exploration.
Under the use of the models, we have created immersive reconstructions in the form of an interactive VR environment. This iVR environment serves as proof of concept of the potential of IA in archaeological analysis. We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatiotemporal data.
Our iVR environment allows users to manipulate lighting conditions, toggle between showing a colour coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor. They use their hands to change the lighting and gain haptic feedback when using buttons inside the virtual space. We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension (Marriott et al., 2018). We chose VR because the ability to interact with the reconstruction in an immersive environment supports a deeper understanding of spatial relationships. This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy (Ragan et al., 2012). The colour coding we use builds upon established practices in Virtual Archaeology to represent reliability in reconstructions (Demetrescu, 2018).
Using methodologies from both fields shows that their respective strengths can be combined. The feature space of the respective topics in IA and Virtual Archaeology is vast, but we did not integrate a large number of potentially valuable features into our software. Our objective was not to provide comprehensive coverage but the demonstration of a number of possibilities as an example.
The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching through them on demand. Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those which not known but can be derived from similar architecture. Changing the angle of incidence of the light source in combination with the miniature to see the relative position helps to gain an idea about lighting, especially inside the different interpretations of the mausoleum’s rotunda.
The general implications of this work are significant for both archaeological methodology and public engagement. It emphasises a number of ways to interact with and analyse archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from. The ability to explore spatial relationships, lighting conditions, and levels of certainty in an immersive environment with embedded analysis information could lead to new research questions. Also, the presentation of data to people with and without domain background with the aspiration to make underlying reasoning understandable will benefit, as well. Scientifically, it emphasises a number of ways to interacting with and analysing archaeological data in terms of representation and metadata.
Due to the character of this work, there are a number of limitations to consider. As for the models we created, as is usual in the domain, they also have parts that are more interpretative than others. The current implementation is focused on a single site and a specific set of interaction techniques. Although multi-modality is important, only haptic feedback and a visual representation of the models are available. Additionally, the use of VR hardware limits the accessibility for some users. While we do offer simple controls that allow to use it without VR hardware, the immersion factor that comes with VR is lost. Also, the considerations we make are a number of potential intersections of the two fields without taking deep considerations of how the design space of the different methods, like interaction design, would have to be applied.
Thus, future work should include a variety of aspects. For one, a unified framework for creating immersive experiences would pose an asset. Such a framework should cover the inclusion and on-demand availability of historical background inside an immersive environment or the sonification of information, both of which have benefits. It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time and date. Another feature to further improve such software would be the simulation of crowds, as seen in the work on the Angkor Wat in Cambodia (Marriott et al., 2018, pp. 346–354).
However, not only software is required. A relevant direction is to assess specific and established workflows from IA and how they can be transferred and used in Virtual Archaeology. It could also be valuable to compare the effectiveness of immersive environments with traditional methods of archaeological data analysis. Lastly, it would be interesting to see the effect of new metaphors for representing inherent information, such as levels of certainty, volumes of objects, sizes, usage, or chronological embedding.
Therefore, there are a great number of opportunities in the future in this intersection, which should be discovered and evaluated together with domain experts to gain further insights into the usage of the Maxentius Mausoleum specifically and cultural heritage objects in general.
7. Conclusion
This work highlights Immersive Analytics’s potential for Virtual Archaeology, particularly for virtual reconstruction accessibility. We briefly discuss the possibilities of engaging representations of reconstructions and interactive representations of hypotheses. We exemplify the concept of integrating IA with our current research of the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding the construction and use of the Maxentius Mausoleum. To tackle the existing debates and derive new research questions, we provide an environment that showcases a few possibilities apart from standard visualisation. We virtualise the mausoleum’s reconstruction models and allow for some manipulation of the environment as well as changing and colourising the models akin to the Extended Matrix by Demetrescu et al. For example, we provide a general impression of how a light source must be positioned to obtain enough light inside the mausoleum, which may help rule out specific window configurations. Rasch’s proposed opening at the top seems to provide enough light to illuminate the whole room, while Johnson’s windows, as they are found at Tor de’ Schiavi, make it challenging to achieve that without additional light sources inside the mausoleum.
In our future research, we want to look into how the workflows from IA can be integrated into digital reconstruction. We want to combine existing modelling workflows for cultural heritage objects and advancements in data representation in immersive environments. We will continue to utilise the Maxentius Mausoleum for that and branch out to other objects, where appropriate. Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.
8. Acknowledgements
We acknowledge funding by DFG, project ID 251654672 – TRR 161. We also acknowledge the support of Antonella Bidlingmaier, who contributed to the creation of the first iterations of the reconstructions of the Maxentius Mausoleum.

### v2

@@
-The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius
-Abstract.
-Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation by, e.
-g., improved data recording technologies and virtual reality devices.
-Immersive environments provide novel ways to present historical events or objects with high visual quality for both, the general public as well as researchers.
-Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics, utilising reconstructions of the mausoleum of the late Roman emperor Maxentius in Rome for argumentation.
-Based on our work, we advocate the value of combining historical and computer science expertise for virtual reconstructions and immersive environments to facilitate deeper understanding and interactive exploration of archaeological data.
-copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JOCCHjournalvolume: xjournalnumber: xarticle: xpublicationmonth: 1ccs: Applied computing Architecture (buildings)ccs: Human-centered computing Visualization application domainsccs: Applied computing Computers in other domainsccs: Human-centered computing Computer supported cooperative workccs: Applied computing Archaeology
-1. Introduction
-Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualisation for work on cultural heritage (Reilly, 1990; Prescott, 1997).
-Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen immense development in its technological capabilities in terms of data acquisition and data representation, as well as several definitions of charters which employ meaningful standards for visualisations and virtual reconstructions (Lopez-Menchero and Grande, 2011; Denard, 2012).
-According to the Seville Charter (Lopez-Menchero and Grande, 2011), Virtual Archaeology concerns itself with finding comprehensive ways of visualising cultural heritage objects.
-The visualisation of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e.
-g., models in virtual reality), since 2D visualisations often fall short of capturing the complexity and spatial relationships inherent in archaeological data.
-Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino (Milgram and Kishino, 1994).
+The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius thanks: DFG, project ID 251654672 – TRR 161
+Abstract
+Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation, for example, by improved data recording technologies and virtual reality devices.
+Immersive environments provide novel ways to present historical events or objects with high visual quality for both the general public and researchers.
+Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics.
+We discuss a selection of features already used by the community and look into how optimizing those can facilitate the discourse and impression of cultural heritage objects.
+As an argumentational basis, we introduce and utilize three digital reconstruction interpretations of the mausoleum of the late Roman emperor Maxentius in Rome, which are based on prior scientific work and a typological embedding.
+Based on our work, we advocate the value of combining historical and computer science expertise to optimize immersive environments for virtual reconstructions to facilitate deeper understanding and interactive exploration of archaeological data.
+Index Terms:
+virtual archaeology; cultural heritage; 3D reconstruction; immersive analytics; virtual reality; Maxentius mausoleum;
+I Introduction
+Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualization for work on cultural heritage [1, 2].
+Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen strong development in its technological capabilities in terms of data acquisition and data representation, as well as definitions of charters that employ meaningful standards for visualizations and virtual reconstructions [3, 4].
+According to the Seville Charter [3], Virtual Archaeology concerns itself with finding comprehensive ways of visualizing cultural heritage objects.
+The visualization of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e.
+g., models in virtual reality) since 2D visualizations often fall short of capturing the complexity and spatial relationships inherent in archaeological data.
+Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino [5].
 Most commonly, immersive technologies are understood to be either virtual reality (VR) as an entirely virtual environment, augmented reality (AR) as an overlay of additional information into the real world, or mixed reality (MR) as a blend between the virtual and real world.
-By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artefacts but also reconstructed objects, background information, or a contextually appropriate environment.
-This leads to a more immersive and interactive exploration of the past.
-Rome, as a focal point of archaeological work in Europe, enjoys a high popularity for such work, leading to an increase in more extensive reconstruction visualisations.
-An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD (Frischer and Massey, 2022, pp.
-23–25).
-However, since such reconstructions target a broader audience, their representations are often suggestive.
-The transparency of their underlying archaeological data origin tends to be selective, and their authenticity and reliability are therefore limited.
-This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction (Liverani, 2021, pp.
-13–14).
-However, to support scientific debates, digital reconstructions can be used to develop innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualise different chronological phases of archaeological sites, and reflect on historical monuments’ construction history and building process.
-A clear indication of different phases, information sources or certainty of validity can, among others, take the form of colour mappings or opacity of the respective components integrated into the visualisation (Bartz et al., 2016; Demetrescu, 2018; Pietroni and Ferdani, 2021).
+By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artifacts but also reconstructed objects, background information, or a contextually appropriate environment.
+Rome, as a focal point of archaeological work in Europe, enjoys high popularity for such work, leading to an increase in more extensive reconstruction visualizations.
+An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD [6, pp.
+23–25].
+The transparency of the underlying archaeological data origin of such public audience-oriented applications tends to be selective, and their authenticity and reliability are therefore limited.
+This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction [7, pp.
+13–14].
+To support scientific debates, digital reconstructions are useful for developing innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualizing different chronological phases of archaeological sites, and reflecting on historical monuments’ construction history and building process.
+A clear indication of different phases, information sources, or certainty of validity can, among others, take the form of color mappings or opacity of the respective components integrated into the visualisation [8, 9, 10].
 To combine the potential of immersive environments and scientific debates, the relatively young field of Immersive Analytics (IA) offers plenty of advancements.
-IA uses engaging and embodied tools to support data understanding (Marriott et al., 2018), explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas.
-Therefore, IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality.
-As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it (Bienroth et al., 2022; Nowke et al., 2013).
-The investigation of which navigation or interaction methods are most appropriate for a specific context (Drogemuller et al., 2018), how effective VR is in teaching contexts (Sari et al., 2023), or the impact of different modalities like audio or haptics (Ruotolo et al., 2013; Venkatesan et al., 2023) are some of its core questions.
-The answers to those questions are as relevant for cultural heritage dissemination and research as they are for other applications.
-They support understanding of findings and to develop hypotheses, which in a reconstruction context means deciding on more meaningful interpretations of the original historical intent and use.
-In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, like annotating specific objects or positions or recording workflows.
-Creating immersive environments that allow people to explore sites, artefacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects, supports bridging the gap between academia and the public.
-Current scientific debates can be communicated by condensing the underlying information and providing a more relatable and captivating experience.
-Since immersive experiences are broadly utilised in the context of dissemination, like museums or teaching, as well as for research (Boboc et al., 2022; Rodriguez-Garcia et al., 2024), and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies (Münster, 2019), we discuss perspectives of how the mentioned potentials of IA can be utilised.
-We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions, and utilise three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig.
+IA uses engaging and embodied tools to support data understanding [11], explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas.
+IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality.
+As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it [12, 13].
+The investigation of which navigation or interaction methods are most appropriate for a specific context [14], how effective VR is in teaching contexts [15], or the impact of different modalities like audio or haptics [16, 17] are some of its core questions.
+In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, such as by investigating how to annotate specific objects or positions or how to collaborate in which contexts.
+Creating immersive environments that allow people to explore sites, artifacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects supports bridging the gap between academia and the public.
+Thus, it is crucial to emphasize the optimization of those supporting features based on the cultural heritage item displayed and the user group towards which the virtual environment is targeted.
+Since immersive experiences are broadly utilized for dissemination, like museums or teaching, as well as for research [18, 19], and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies [20], we discuss perspectives of how the mentioned potentials can be utilized.
+We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions.
+We use three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig.
 1).
-Those reconstructions are based on ongoing scientific debates on what the mausoleum might have looked like in its prime, and we provide those in addition to software to inspect them in VR.
-Thus, our contribution is a perspective on where those two research directions can be merged to further discourse and show, by example, some of the possible concepts and techniques.
-Maxentius A                          Maxentius B                        Maxentius C
-Figure 1.
-This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work.
-The first model (Maxentius A) closely follows Rasch’s reconstruction proposal (Rasch, 1984).
-The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi (Johnson, 2009).
+They are based on ongoing scientific debates about what the mausoleum might have looked like in its prime.
+We provide these reconstruction models in addition to software to inspect them in VR.
+Thus, our contribution is comprised of a perspective on how Immersive Analytics research can help to assist archaeological discourse and the three digital reconstructions of the Maxentius mausoleum that we created.
+Furthermore, we provide an example software to demo some of the possible concepts and techniques.
+Figure 1: This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work.
+The first model (Maxentius A) closely follows Rasch’s reconstruction proposal [21].
+The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi [22].
 The third model (Maxentius C) tries to merge the dome and the windows of Tor de’ Schiavi to the proportions of the Maxentius Mausoleum.
-[…]
-Renderings of all three digital reconstruction alternatives we propose next to each other with a focus on the different roofs, which pose the main difference between those models.
-2. Background
-Given the interdisciplinary nature of the topic, this section will introduce key vocabulary that is contextually relevant.
-In the work of Lopez-Menchero et al.
-(Lopez-Menchero and Grande, 2011), virtual reconstruction is defined as using a virtual model to recover a building or object made by humans visually.
-Here, we differentiate between digital reconstructions, virtual reconstructions, and immersive reconstructions to indicate the different contexts.
-Digital reconstruction is the process of digitally reconstructing an object through digital (computational) tools.
-Virtual reconstruction is the representation of a cultural heritage object in a virtual space to explore and interact with it.
-It is an application of digital reconstruction.
-Immersive reconstruction is the representation of a cultural heritage object in an immersive environment, that is, an environment along Milgram’s reality-virtuality continuum.
-It can be virtual, augmented, or mixed and can include additional techniques, such as data physicalisation, to provide an immersive experience.
-One term relevant for reconstructions is anastylosis as part of virtual restoration, which describes the process of using in-situ findings as a reference, where in-situ means that they are found at their original cultural heritage site or, more specifically, directly “where they belong”.
-Those findings are then placed at their original position in the reconstruction and utilised to infer further missing components.
-This workflow is thoroughly described and used as part of the Extended Matrix by Demetrescu et al.
-(Demetrescu, 2018), who use colour coding to represent the reliability between different components of a reconstruction from still existing in-situ components which reside at their original position to purely hypothetical parts.
-As we will discuss later in this work, virtual and immersive reconstructions often consist not only of the cultural heritage object itself but also of an artificial environment which, to some degree, resembles the environment where the object was formerly to be found.
-This concept adheres to the idea of situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making (Marriott et al., 2018, p.
-185).
+II Background
+The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching, and even research in archaeological contexts has grown [18, 19, 6, 23, 24].
+As we focus on reconstructions in this work and its interdisciplinary nature, it’s important to clarify the different types of reconstructions in the digital realm.
+Digital reconstruction involves creating objects using computational tools, while virtual reconstruction is the use of virtual models to recover a building or man-made object visually [3].
+Immersive reconstruction extends this further by embedding objects in environments along Milgram’s reality-virtuality continuum [5], such as AR or VR, to enhance user experience through interaction and context.
+As is the nature of a reconstruction, it is not granted that all pieces of a monument, for instance, are known.
+Researchers have to complement the known parts with modeled ones, whose form and structure derive from different sources, like paintings, folklore, derivation from existing pieces, and so on.
+These different sources need to be pointed out, as Demetrescu describes in his work [9].
+The sources of the reconstruction components, as well as the surrounding area and information on the object’s purpose or its material, are all data that can be transported via various means, especially so with the help of immersive technologies.
+Immersive technologies can support reasoning [25] by embedding data into the physical world [26], the recall of data [27], or spatial judgment accuracy [28].
+This is also relevant in Virtual Archaeology, as such environments have proven to create a heightened sense of presence in its’ context, as already shown by Kyrlitsias, Christofi, Michael-Grigoriou, Banakou, and Ioannou [29].
+Presence, as a form of psychological immersion, can also be called engagement [30], and is an important concept, as there is evidence that task performance and recall are boosted through heightened engagement [31].
+The research direction that concerns itself with investigating how to exploit those engaging tools to increase data understanding is called Immersive Analytics [11].
+In their book on Immersive Analytics, Marriott et al.
+describe a number of relevant aspects apart from those already discussed.
+One of those is Situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making [32].
 For example, it might overlay an existing object with analysis information about it and is, thus, tightly linked to AR applications.
-Embodiment relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point and click interaction as it is typical for desktop PCs (Marriott et al., 2018, p.
-3, p.
-101).
+Another is the Embodiment, which relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point-and-click interaction as it is typical for desktop PCs [30].
 That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell.
-The psychological immersion, also called engagement (Marriott et al., 2018, p.
-10), is another important concept, as there is evidence that task performance and recall are boosted through heightened engagement (Makowski et al., 2017).
-The last concept to mention is collaborative immersive analytics.
-In IA, many problems addressed require more than one person to solve (Marriott et al., 2018, p.
-221), leading to multiple people working in the same virtual environment.
-This includes the collaboration of multiple people from different physical locations and even at various points in time.
-3. Immersive Analytics for Virtual Archaeology
-The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching and even research in archaeological contexts has grown (Boboc et al., 2022; Rodriguez-Garcia et al., 2024).
-As we focus on reconstructions in this work, we consider digital reconstruction models as well as the data recordings, which lead to those being the data we work with and seek to understand.
-Consequently, the general intersection with IA is clear: We use current technologies to present our data, and we want to optimise the experience to further a user’s ability to understand the reconstructions.
-To better understand data, some level of engagement is desirable, as there is evidence for an increased recall of data (Mark et al., 2003).
-This can be achieved by using immersive VR environments, which are popular in different outreach-related domains (Frischer and Massey, 2022; Schofield et al., 2018; Kyrlitsias et al., 2020).
-Such environments have proven to create a heightened sense of presence, as already argued by Kyrlitsias et al.
-(Kyrlitsias et al., 2020), which matches findings from IA.
-For example, Ragan et al.
-argue that the combination of s3D and head-tracking helps spatial judgment accuracy (Ragan et al., 2012).
-The analysis they conducted in their study looked into the different effects of stereovision, head-tracking and bodily movement by tasking their participants with context-related tasks, in which the participant’s error rate was almost halved when using stereovision and head-tracking.
-Nowadays, such analysis can easily be done for cultural heritage dissemination environments, as many modern head-mounted display devices have integrated telemetry hardware like eye-tracking, and all of them have either internal or external head-tracking.
-Museums, for instance, could collect data to see which points of interest in their exhibit attract visitors.
-In general, an investigation could be conducted on how people perceive any given environment and whether the scene is overloaded or achieves the desired recall of the presented information.
-AR is also already seeing applications in virtual archaeology  (Tscheu and Buhalis, 2016).
-Such environments create a heightened sense of presence, which matches findings from IA that situated analytics supports reasoning (ElSayed et al., 2015) by embedding data into the physical world (Willett et al., 2017).
-Several papers on situated analytics cover pipelines for creating corresponding environments (Kalkofen et al., 2007; Willett et al., 2017; Fleck et al., 2023), leading to the conclusion that these pipelines can be used for virtual reconstruction models as well.
-One widespread use for AR is mapping digital information onto a physical referent, an approach described by the conceptual model by Willett et al.
-(Willett et al., 2017).
-Given such a physical referent is a heritage site or a part of it, reconstructed components can be acquired from an existing data set like digital twins, digital reconstructions or even just raw point clouds, and the actual physical object can be supplemented by a virtual representation visible in the AR system.
-For example, a proximity-based magical lense visualisation (Lee et al., 2023) could be used, where supplement information for an object is only shown in an embedded fashion given a certain proximity of the user to the object.
-Other known or inferred information can be overlayed apart from reconstructed components.
-This could be, for example, the approximate age of an illustration or an estimated purpose of an object, which could be calculated in the background, or the “Logical World”, as Willett et al.
-call it.
-Another information that could be provided that way is the different types of archaeological evidence as described by Demetrescu et al.
-(Demetrescu, 2018), from actual residuals of the degraded objects to inferences from typologically similar objects or lore.
-Those types of evidence can be mapped to different levels of certainty, from well-preserved to purely hypothetical, and should be highlighted appropriately in reconstruction visualisations (Demetrescu, 2018; Schäfer, 2018; Pietroni and Ferdani, 2021).
-While visualisation is essential, current immersive environment devices offer a wide variety of possibilities to utilise different modalities for input and output.
-One of which is haptics, which supports sense-making for priorly unseen data (Taher et al., 2017).
-Encoding information like material density or even the age of different objects into the intensity of haptic feedback are possible channels for additional information.
-Furthermore, data physicalisation, for instance, in the form of 3D prints, is already being used to provide a deeper understanding of existing information (Klein et al., 2022), which gives users the possibility of actually touching their data.
-Such physicalisations could, in turn, be used as a way of interacting with reconstructions.
-For example, an AR device could recognise the region a user is pointing at to provide information about it.
-Another multi-modal channel for in- and output is sound, and there are various ways of using sound to map data (sonification) (Madhyastha and Reed, 1995).
-Having access to realistic sound models can already help to interpret which purpose a specific room might have served (McBride, 2014).
-Since ambient sound can reduce distraction and improve presence (Kern and Ellermeier, 2020), the identification and application of contextually appropriate ambient sounds which are based on a user position (in the digital world) or activities, as well as attention control through specifically designed noises can be used to improve dissemination quality or aid validity assessment.
-Further intriguing aspects of IA environments are the possibilities that arise from development kits such as Unity (Unity Technologies, 2023) or Unreal Engine (Epic Games, 2024), which come with working physics engines and multi-user capabilities.
-Through physics engines, concepts like lighting can be realistically simulated to allow for the investigation of their influence on reconstructions.
-Through multi-user capabilities, multiple people who do not necessarily have to be at the same place (distributed collaboration) can look at reconstructions and interact with them, which augments the understanding of users (Hill, 1982; Mark et al., 2003).
-By integrating user interfaces, either simple VR versions of desktop-like interfaces or more natural ones (Wigdor, 2011), users can furthermore be given the power to manipulate objects or the environment as a whole directly.
-Definitions and a more generalised approach towards interaction systems can be found in IA-related work (Büschel et al., 2018).
-Such interaction systems can be utilised and combined to reasonable manipulation methods of a cultural heritage object, such as interactively changing the lighting conditions, adding annotations or filters for components, or reassembling pieces of the reconstructions in a playful manner (Houska et al., 2024).
-With a selection of those considerations, we will further discuss the example of the Maxentius Mausoleum and how these considerations can be applied to an environment for it.
-4. Maxentius Mausoleum
-Figure 2. The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
-[…]
-A photo we took of the Maxentius Mausoleum in 2024. Visible are the remains of the crypt from the outside as well as the farmhouse.
+III Maxentius Mausoleum
+Figure 2: The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
 To provide insight into the significance of our archaeological example, we briefly discuss its historical and typological embedding and connection to the reign of Maxentius.
@@
 306-312) gave Rome again the status of a permanent imperial residence.
-Maxentius styled himself as the conservator of Rome (conservator urbis suae) and added several large buildings such as the Maxentius Basilica and the temple of Venus and Rome, thermal baths and an extension of the imperial palace on the Palatine Hill to the architectural ensemble of the city (Leppin and Ziemssen, 2007, pp.
-12-13, 51, 74-82, 87-95).
-On the city’s outskirts, at the third milestone of the Via Appia leading southwards, Maxentius built a residential complex, which encompassed a villa suburbana with an attached well-preserved circus and a monumental mausoleum directed towards the Via Appia.
-The mausoleum, which forms the core of the complex, is surrounded by a slightly offset portico of 99.50 by 85 meters, of which substantial remains (as high as 11 meters) can still be seen (Rasch, 1984, p.
-14) (see Fig.
-2).
-However, the preserved remains of the mausoleum give the modern visitor only a limited idea of its former importance and structure.
-Just the foundations of the original building remain, overbuilt partly by a farmhouse of the 18th century (Rasch, 1984, pp.
-5–6).
-Inside the building, an ambulatory crypt around an overarching pillar with niches inside is well-preserved.
-The layout of the building includes a round podium with a diameter of 33.25 meters.
-Adjacent is a front porch partly overbuilt by a farmhouse with a length of 17.00 meters and a width of 22.68 meters.
-Covered by the farmhouse, a vaulted cellar of the front porch with a diameter of 9.10 by 16.33 meters remains.
-The still-existing crypt has a diameter of 23.54 meters, a central column 9.70 meters wide, and seven niches (Johnson, 2009, pp.
-86–90).
-The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD (Johnson, 2009, pp.
-92–93).
+Maxentius styled himself as the conservator of Rome (conservator urbis suae) and added several large buildings such as the Maxentius Basilica and the temple of Venus and Rome, thermal baths, and an extension of the imperial palace on the Palatine Hill to the architectural ensemble of the city [33, pp.
+12-13, 51, 74-82, 87-95].
+[…]
+The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD [22, pp.
+92–93].
 Its few remaining traces provide an excellent case study for the application of virtual archaeology.
 Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon.
-From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modelled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey (Rasch, 1993, p.
-84).
-However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) (Rasch, 1993).
+From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modeled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey [34, p.
+84].
+However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) [34].
 This monument is located, like the Maxentius Mausoleum, on the outskirts of Rome, close to an imperial villa suburbana.
@@
 These different perspectives, into which Maxentius’ Mausoleum can be classified, result in different conceivable models for its reconstruction, which will be discussed below.
-5. Immersive Analytics for the Maxentius Mausoleum
+IV Immersive Analytics for the Maxentius Mausoleum
 Regarding the Maxentius Mausoleum, we chose Rasch’s work as a foundation for a first reconstruction (Maxentius A), which closely follows his suggested measurements.
-The second model (Maxentius B) assumes a closed dome following Johnson’s proposal (Johnson, 2009, pp.
-91-92) and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi (Rasch, 1993, p.
-28).
-The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi (Rasch, 1993, p.
-68) and the windows with a larger diameter.
-Figure 3.
-Cross-faded version of Maxentius A.
-Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
-[…]
-A transparent version of a digital reconstruction based on Rasch’s work is put on a photogrammetry-based digital twin of the current state of the mausoleum.
-This model visualises, as suggested by Johnson, a version of the Maxentius Mausoleum modelled closely on Tor de’ Schiavi (Johnson, 2009, p.
-92).
+The second model (Maxentius B) assumes a closed dome following Johnson’s proposal [22, pp.
+91-92] and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi [34, p.
+28].
+The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi [34, p.
+68] and the windows with a larger diameter.
+This model visualizes, as suggested by Johnson, a version of the Maxentius Mausoleum modeled closely on Tor de’ Schiavi [22, p.
+92].
 Consequently, model B can be considered a merger of models A and C.
@@
 1), which mainly differ in the dome’s design and allow us to discuss which of the hypotheses about the part of the mausoleum is most plausible.
-We created the reconstructions with the 3D modelling software Blender.
+We created the reconstructions with the 3D modeling software Blender.
 The components of the models consist of combined simple geometric forms, which were manipulated through manual node shifting to obtain necessary details.
@@
 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
-Figure 4. The software we provide together with our models. The panel, the miniature, and the actual visualisation can be seen.
-[…]
-An image from a user’s perspective, where the user sees an in-scale reconstruction of the Maxentius Mausoleum.
-What the user also sees is a miniature of the mausoleum reconstruction on a cube as well as a small panel with buttons with the names Maxentius A, B, and C, referring to the three available reconstructions.
+Figure 3: Cross-faded version of Maxentius A.
+Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
 We provide software (see Fig.
-4) that presents the three hypothesised models of the Maxentius mausoleum.
-The software provides a Virtual Reality environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index.
-Inside the immersive environment, the user will start in front of the stairway of Maxentius A.
-The user can move around, using the continuous movement metaphor and can snap-turn or move around freely, depending on the physical space available around them.
-When starting the software, in front of the user is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a colouring akin to the extended matrix by Demetrescu et al.
-The buttons can be touched by the user with their hands.
-When the user touches one of the buttons, the presented model will switch, and the user will get haptic feedback.
+4) that presents the three hypothesized models of the Maxentius mausoleum.
+The software provides a VR environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index.
+We chose VR as it represents an s3D environment with head-tracking, which allows for improved spatial judgment.
+Inside the immersive environment, the user will start in front of the Maxentius A stairway.
+The user can move around, using the continuous movement metaphor, and can snap-turn or move around freely, depending on the physical space available around them.
+Figure 4: The software we provide together with one of our models. The panel, the miniature, and the actual visualization can be seen.
+There is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a coloring akin to the extended matrix by Demetrescu by touching those buttons with their hands.
+Switching between models is especially relevant in a research context for comparing several different hypotheses directly while still in the iVR environment.
+Consequently, in dialogue with domain experts, compiling ways to compare the reconstruction models intuitively or conveniently helps to improve cognitive affordance.
 To the right of the panel is a miniature of the model.
-Users can hold their virtual hands inside it and grab it with a simple grabbing motion.
-Upon grabbing, the user can control with their wrist the position of the sun and thus the light’s angle of incidence on the model.
-There is a small yellow sphere indicating the relative position of the light source, which imitates the sun so that users do not have to search for the real sun’s position.
-Once the user reaches for the miniature, a vibration of the controller indicates that they can now grab to rotate the light source.
-Once the grabbing motion is performed, a slightly more intense and more prolonged vibration pattern means that the wrist’s rotation now manipulates the light source’s angle.
-Furthermore, users can place small squares to leave small randomly coloured markers at points of interest behind.
-Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need of additional hardware.
-We provide the software via a Zenodo upload at (Kerle-Malcharek, 2025).
-6. Discussion
-Our exploration of IA applications in Virtual Archaeology argues for a significant potential for enhancing data understanding and analysis in archaeological reconstructions.
-While there exist works which advocate for the usefulness of immersive environments in Virtual Archaeology (Kyrlitsias et al., 2020), many works focus on dissemination or reconstruction.
-Therefore, this work yields significance for the field by presenting general considerations of how research can benefit from immersive technologies.
-For our considerations, we utilise the Maxentius Mausoleum in Rome as a case study in the form of three digital reconstructions.
-The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea, which we also present, and are based on archaeological evidence and scholarly interpretations.
-We did this since we wanted to use reconstructions based on scientific evidence and valid assumptions to provide a sound foundation for our exploration.
-Under the use of the models, we have created immersive reconstructions in the form of an interactive VR environment.
+By holding their virtual hands inside the miniature and performing a grabbing motion, a vibration of the controller indicates interaction, and they can control the light’s angle of incidence on the model.
+There is a small yellow sphere indicating the relative position of the light source at the miniature.
+This small sphere is intended to serve as a visualization to ease the intuition of where the light comes from.
+This idiom is supposed to help when analyzing how and where sunlight comes from for the mausoleum.
+For our mausoleum scenario, it would make sense to look into relevant ways to interact with light sources.
+An optimized light interaction system could lead to a more fluent experience, which in turn may help experts justify the different shapes of the dome and the light conditions inside.
+Furthermore, users can place small squares to leave small randomly colored markers at points of interest behind.
+In general, different types of markers could be developed.
+For example, users could be enabled to colorize specific areas of the wall to mark the locations of the numerous drawings on the walls, which were added long after the mausoleum was originally built.
+The type of marker, color coding, or maybe even an acoustic signal could then again be investigated for its effectiveness in finding the different spots.
+Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need for additional hardware.
+We provide the software via a Zenodo upload at [35].
+V Discussion
+With this work, we present three digital reconstructions of the Maxentius Mausoleum in Rome and utilize them to assess the value of IA for Virtual Archaeology.
+The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea and are based on archaeological evidence and scholarly interpretations.
+Such scientific evidence provides us with a sound foundation for our discussion in which we want to advocate for the use of immersive technologies more prominently for research purposes.
+While there exist works that point out the usefulness of immersive environments in Virtual Archaeology [29], they tend to focus on dissemination or reconstruction and not on how to improve the perception of a specific aspect, like the comparison of hypotheses.
+Using the three reconstruction models, we created immersive reconstructions in the form of an interactive VR (iVR) environment.
 This iVR environment serves as proof of concept of the potential of IA in archaeological analysis.
-We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatiotemporal data.
-Our iVR environment allows users to manipulate lighting conditions, toggle between showing a colour coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor.
+We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatio-temporal data.
+Our iVR environment allows users to manipulate lighting conditions, toggle between showing a color coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor.
 They use their hands to change the lighting and gain haptic feedback when using buttons inside the virtual space.
-We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension (Marriott et al., 2018).
+We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension [11].
 We chose VR because the ability to interact with the reconstruction in an immersive environment supports a deeper understanding of spatial relationships.
-This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy (Ragan et al., 2012).
-The colour coding we use builds upon established practices in Virtual Archaeology to represent reliability in reconstructions (Demetrescu, 2018).
-Using methodologies from both fields shows that their respective strengths can be combined.
-The feature space of the respective topics in IA and Virtual Archaeology is vast, but we did not integrate a large number of potentially valuable features into our software.
-Our objective was not to provide comprehensive coverage but the demonstration of a number of possibilities as an example.
+This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy [28].
+The color coding we use builds upon established practices in Virtual Archaeology to represent reliability in reconstructions [9].
+As the feature space of IA topics and Virtual Archaeology is vast, we decided against integrating an exhaustive number of potential features into our software.
+Our objective was not to provide comprehensive coverage but to demonstrate a number of possibilities as an example.
 The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching through them on demand.
-Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those which not known but can be derived from similar architecture.
-Changing the angle of incidence of the light source in combination with the miniature to see the relative position helps to gain an idea about lighting, especially inside the different interpretations of the mausoleum’s rotunda.
+Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those that are not known but can be derived from similar architecture.
+Through the miniature, users can see how the light source’s position changes relative to the reconstruction.
 The general implications of this work are significant for both archaeological methodology and public engagement.
-It emphasises a number of ways to interact with and analyse archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from.
+It emphasizes a number of ways to interact with and analyze archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from.
 The ability to explore spatial relationships, lighting conditions, and levels of certainty in an immersive environment with embedded analysis information could lead to new research questions.
-Also, the presentation of data to people with and without domain background with the aspiration to make underlying reasoning understandable will benefit, as well.
-Scientifically, it emphasises a number of ways to interacting with and analysing archaeological data in terms of representation and metadata.
 Due to the character of this work, there are a number of limitations to consider.
@@
 Additionally, the use of VR hardware limits the accessibility for some users.
-While we do offer simple controls that allow to use it without VR hardware, the immersion factor that comes with VR is lost.
+While we do offer simple controls that allow the software without VR hardware, the immersion factor that comes with VR is lost.
 Also, the considerations we make are a number of potential intersections of the two fields without taking deep considerations of how the design space of the different methods, like interaction design, would have to be applied.
@@
 It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time and date.
-Another feature to further improve such software would be the simulation of crowds, as seen in the work on the Angkor Wat in Cambodia (Marriott et al., 2018, pp.
-346–354).
 However, not only software is required.
@@
 Therefore, there are a great number of opportunities in the future in this intersection, which should be discovered and evaluated together with domain experts to gain further insights into the usage of the Maxentius Mausoleum specifically and cultural heritage objects in general.
-7. Conclusion
-This work highlights Immersive Analytics’s potential for Virtual Archaeology, particularly for virtual reconstruction accessibility.
-We briefly discuss the possibilities of engaging representations of reconstructions and interactive representations of hypotheses.
-We exemplify the concept of integrating IA with our current research of the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding the construction and use of the Maxentius Mausoleum.
-To tackle the existing debates and derive new research questions, we provide an environment that showcases a few possibilities apart from standard visualisation.
-We virtualise the mausoleum’s reconstruction models and allow for some manipulation of the environment as well as changing and colourising the models akin to the Extended Matrix by Demetrescu et al.
-For example, we provide a general impression of how a light source must be positioned to obtain enough light inside the mausoleum, which may help rule out specific window configurations.
-Rasch’s proposed opening at the top seems to provide enough light to illuminate the whole room, while Johnson’s windows, as they are found at Tor de’ Schiavi, make it challenging to achieve that without additional light sources inside the mausoleum.
+VI Conclusion
+In this work, we present three reconstruction hypotheses of the Maxentius Mausoleum and the literature and typology that inspired them.
+In addition, we have created an immersive environment in which those reconstructions can be viewed and explored.
+The environment comes with a small number of integrated interaction and navigation possibilities to provide not only the models but also the means to explore them.
+Our second contribution is the exploration of the potential for Virtual Archaeology that Immersive Analytics as a research direction yields.
+We advocate the shift from ever-better representations, recording, and storage of cultural heritage items to tailored immersive experiences that improve accessibility and insight into existing data.
+The direction is to look at how virtual archaeology studies can be optimized by identifying the specific requirements of a given problem, like comparing different monument reconstruction interpretations and finding out how to support that.
+For argumentation, we outline the possibilities of engaging representations of reconstructions and interactive representations of hypotheses.
+We exemplify the concept of integrating IA with our current research on the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding its construction and use.
 In our future research, we want to look into how the workflows from IA can be integrated into digital reconstruction.
-We want to combine existing modelling workflows for cultural heritage objects and advancements in data representation in immersive environments.
-We will continue to utilise the Maxentius Mausoleum for that and branch out to other objects, where appropriate.
+We like to combine existing modeling workflows for cultural heritage objects and advancements in data representation in immersive environments.
+We will continue to utilize the Maxentius Mausoleum for that and branch out to other objects, where appropriate.
 Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.
-8. Acknowledgements
-We acknowledge funding by DFG, project ID 251654672 – TRR 161.
-We also acknowledge the support of Antonella Bidlingmaier, who contributed to the creation of the first iterations of the reconstructions of the Maxentius Mausoleum.
+Acknowledgment
+We acknowledge funding by DFG, project ID 251654672–TRR 161 as well as the support of Antonella Bidlingmaier, who contributed to the creation of the first iterations of the reconstructions of the Maxentius Mausoleum.

### v3

@@
 Abstract
-Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation, for example, by improved data recording technologies and virtual reality devices.
+Virtual archaeology has significantly evolved over the last few decades through advancements in data acquisition and representation; for example, by improved data recording technologies and virtual reality devices.
 Immersive environments provide novel ways to present historical events or objects with high visual quality for both the general public and researchers.
-Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics.
-We discuss a selection of features already used by the community and look into how optimizing those can facilitate the discourse and impression of cultural heritage objects.
-As an argumentational basis, we introduce and utilize three digital reconstruction interpretations of the mausoleum of the late Roman emperor Maxentius in Rome, which are based on prior scientific work and a typological embedding.
-Based on our work, we advocate the value of combining historical and computer science expertise to optimize immersive environments for virtual reconstructions to facilitate deeper understanding and interactive exploration of archaeological data.
-Index Terms:
-virtual archaeology; cultural heritage; 3D reconstruction; immersive analytics; virtual reality; Maxentius mausoleum;
+Here, we examine how the emerging field of immersive analytics can contribute to enhancing the understanding and exploration of archaeological data, and we explore the junction of virtual archaeology and immersive analytics.
+We discuss a selection of features already used by the community and examine how optimizing these can facilitate the discourse on cultural heritage objects.
+As a basis for discussion, we introduce and utilize three digital reconstruction interpretations of the mausoleum of the late Roman Emperor Maxentius in Rome, which are based on prior scientific work and a typological framework.
+Based on our work, we advocate for the value of combining historical and computer science expertise to optimize immersive environments for virtual reconstructions, thereby facilitating a deeper understanding and interactive exploration of archaeological data.
 I Introduction
-Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualization for work on cultural heritage [1, 2].
-Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen strong development in its technological capabilities in terms of data acquisition and data representation, as well as definitions of charters that employ meaningful standards for visualizations and virtual reconstructions [3, 4].
+Terms such as virtual archaeology, virtual restoration, and virtual reconstruction date back to the 1990s, when researchers investigated the potential of virtualization for work on cultural heritage [1, 2].
+Over the years, virtual archaeology has evolved into an entire field of research, with significant advancements in its technological capabilities in terms of data acquisition and data representation, as well as the development of charters that employ meaningful principles for visualizations and virtual reconstructions [3, 4].
 According to the Seville Charter [3], virtual archaeology concerns itself with finding comprehensive ways of visualizing cultural heritage objects.
@@
 Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino [5].
-Most commonly, immersive technologies are understood to be either virtual reality (VR) as an entirely virtual environment, augmented reality (AR) as an overlay of additional information into the real world, or mixed reality (MR) as a blend between the virtual and real world.
+Most commonly, immersive technologies are understood to be either virtual reality (VR), which is an entirely virtual environment; augmented reality (AR), which is an overlay of additional information into the real world; or mixed reality (MR), which is a blend between the virtual and real worlds.
 By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artifacts but also reconstructed objects, background information, or a contextually appropriate environment.
@@
 The transparency of the underlying archaeological data origin of such public audience-oriented applications tends to be selective, and their authenticity and reliability are therefore limited.
-This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction [7, pp.
+This is only natural because the goal of an immersive environment targeted at the broader public is not to discuss different scholarly opinions regarding the reconstruction [7, pp.
 13–14].
-To support scientific debates, digital reconstructions are useful for developing innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualizing different chronological phases of archaeological sites, and reflecting on historical monuments’ construction history and building process.
-A clear indication of different phases, information sources, or certainty of validity can, among others, take the form of color mappings or opacity of the respective components integrated into the visualisation [8, 9, 10].
-To combine the potential of immersive environments and scientific debates, the relatively young field of Immersive Analytics (IA) offers plenty of advancements.
-IA uses engaging and embodied tools to support data understanding [11], explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas.
+To support scientific debates, digital reconstructions are useful for developing innovative questions about the historical perception and sociopolitical context of monuments and objects, visualizing different chronological phases of archaeological sites, and reflecting on the construction history and building process of historical monuments.
+A clear indication of different phases, information sources, or certainty of validity can, among others, take the form of color mappings or opacity of the respective components integrated into the visualization [8, 9, 10].
+To combine the potential of immersive environments with scientific debates, the relatively young field of immersive analytics (IA) offers numerous advancements.
+IA utilizes engaging and embodied tools to support data understanding [11], explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored for specific application areas.
 IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality.
 As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it [12, 13].
-The investigation of which navigation or interaction methods are most appropriate for a specific context [14], how effective VR is in teaching contexts [15], or the impact of different modalities like audio or haptics [16, 17] are some of its core questions.
+The investigation of which navigation or interaction methods are most appropriate for a specific context [14], how effective VR is in teaching contexts [15], or the impact of different modalities such as audio or haptics [16, 17] are only some of its research topics.
 In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, such as by investigating how to annotate specific objects or positions or how to collaborate in which contexts.
-Creating immersive environments that allow people to explore sites, artifacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects supports bridging the gap between academia and the public.
+Creating immersive environments that allow people to virtually explore sites, artifacts, and reconstructed objects, add annotations and landmarks for idea communication, and obtain more than mere visual impressions of cultural heritage objects supports bridging the gap between academia and the public.
 Thus, it is crucial to emphasize the optimization of those supporting features based on the cultural heritage item displayed and the user group towards which the virtual environment is targeted.
-Since immersive experiences are broadly utilized for dissemination, like museums or teaching, as well as for research [18, 19], and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies [20], we discuss perspectives of how the mentioned potentials can be utilized.
-We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions.
+Since immersive experiences are widely utilized for dissemination, such as in museums or teaching, as well as for research [18, 19], and evidence suggests that discourse in virtual archaeology is primarily driven by advancing technologies [20], we discuss perspectives on how the mentioned potentials can be utilized.
+We consider how immersive analytics methods can be used for virtual archaeology, specifically for virtual reconstructions.
 We use three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig.
@@
 They are based on ongoing scientific debates about what the mausoleum might have looked like in its prime.
-We provide these reconstruction models in addition to software to inspect them in VR.
-Thus, our contribution is comprised of a perspective on how Immersive Analytics research can help to assist archaeological discourse and the three digital reconstructions of the Maxentius mausoleum that we created.
-Furthermore, we provide an example software to demo some of the possible concepts and techniques.
-Figure 1: This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work.
+We provide these reconstruction models, along with software, to inspect them in VR.
+Thus, our contribution consists of a perspective on how immersive analytics research can assist archaeological discourse, as well as the three digital reconstructions of the Maxentius mausoleum that we created.
+Furthermore, we provide an example software to demonstrate some of the possible concepts and techniques.
+Figure 1: This image illustrates the three proposed models of the Mausoleum of Maxentius discussed in this work.
 The first model (Maxentius A) closely follows Rasch’s reconstruction proposal [21].
-The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi [22].
-The third model (Maxentius C) tries to merge the dome and the windows of Tor de’ Schiavi to the proportions of the Maxentius Mausoleum.
+The second model (Maxentius B) is based on Rasch’s model but features a closed dome and seven windows in the clerestory, modeled after the Tor de’ Schiavi [22].
+The third model (Maxentius C) attempts to merge the dome and windows of Tor de’ Schiavi with the proportions of the Maxentius Mausoleum.
 II Background
@@
 As we focus on reconstructions in this work and its interdisciplinary nature, it’s important to clarify the different types of reconstructions in the digital realm.
-Digital reconstruction involves creating objects using computational tools, while virtual reconstruction is the use of virtual models to recover a building or man-made object visually [3].
+Digital reconstruction involves creating objects using computational tools, while virtual reconstruction is the use of virtual models to recover a building or manmade object visually [3].
 Immersive reconstruction extends this further by embedding objects in environments along Milgram’s reality-virtuality continuum [5], such as AR or VR, to enhance user experience through interaction and context.
-As is the nature of a reconstruction, it is not granted that all pieces of a monument, for instance, are known.
-Researchers have to complement the known parts with modeled ones, whose form and structure derive from different sources, like paintings, folklore, derivation from existing pieces, and so on.
-These different sources need to be pointed out, as Demetrescu describes in his work [9].
+As is the nature of a reconstruction, it is possible that not all pieces of a monument, for instance, are known.
+Researchers must complement the known parts with modeled ones, whose form and structure derive from various sources, such as paintings, folklore, and derivations from existing pieces.
+These different sources need to be indicated, as Demetrescu describes in his work [9].
 The sources of the reconstruction components, as well as the surrounding area and information on the object’s purpose or its material, are all data that can be transported via various means, especially so with the help of immersive technologies.
 Immersive technologies can support reasoning [25] by embedding data into the physical world [26], the recall of data [27], or spatial judgment accuracy [28].
-This is also relevant in Virtual Archaeology, as such environments have proven to create a heightened sense of presence in its’ context, as already shown by Kyrlitsias, Christofi, Michael-Grigoriou, Banakou, and Ioannou [29].
+This is also relevant in virtual archaeology, as such environments have proven to create a heightened sense of presence, as already shown by Kyrlitsias, Christofi, Michael-Grigoriou, Banakou, and Ioannou [29].
 Presence, as a form of psychological immersion, can also be called engagement [30], and is an important concept, as there is evidence that task performance and recall are boosted through heightened engagement [31].
-The research direction that concerns itself with investigating how to exploit those engaging tools to increase data understanding is called Immersive Analytics [11].
+The research direction that concerns itself with investigating how to exploit tools which engage the user to increase data understanding is called immersive analytics [11].
 In their book on immersive analytics, Marriott et al.
-describe a number of relevant aspects apart from those already discussed.
-One of those is Situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making [32].
-For example, it might overlay an existing object with analysis information about it and is, thus, tightly linked to AR applications.
-Another is the Embodiment, which relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point-and-click interaction as it is typical for desktop PCs [30].
+describe several additional relevant aspects apart from those already discussed.
+One of these is situated analytics, which places data representations in relation to relevant objects, places, and persons to enhance understanding and decision-making [32].
+For example, it might overlay an existing object with analysis information about it and is thus tightly linked to AR applications.
+Another is the embodiment, which relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance, rather than the point-and-click interaction typical of desktop PCs [30].
 That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell.
 III Maxentius Mausoleum
-Figure 2: The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
+Figure 2: The current state of the Maxentius mausoleum. Only the crypt still remains. The front porch is partly missing and has partly been used to build a farmhouse.
 To provide insight into the significance of our archaeological example, we briefly discuss its historical and typological embedding and connection to the reign of Maxentius.
@@
 Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon.
-From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modeled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey [34, p.
+From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a reconstruction of the building modeled on the Pantheon with its specific architectural components: a podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey [34, p.
 84].
@@
 Thus, we have three reconstruction models of the same object (see Fig.
-1), which mainly differ in the dome’s design and allow us to discuss which of the hypotheses about the part of the mausoleum is most plausible.
+1), which mainly differ in the design of the dome, allowing us to discuss which of the hypotheses about that part of the mausoleum is most plausible.
 We created the reconstructions with the 3D modeling software Blender.
@@
 For each sub-component, once constructed from primitives, the measurements for those components were linearly scaled and applied.
-Thanks to the precise measurements we get from Rasch’s work, the models provide correct proportions, which helps us to obtain a fitting impression of the mausoleum’s former glory.
+Thanks to the precise measurements obtained from Rasch’s work, the models provide correct proportions, which help us obtain a fitting impression of the mausoleum’s former glory.
 In Fig.
 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
-Figure 3: Cross-faded version of Maxentius A.
-Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
+Figure 3: Cross-faded version of Maxentius A. The different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
 We provide software (see Fig.
-4) that presents the three hypothesized models of the Maxentius mausoleum.
+4) that presents the three hypothesized models of the Mausoleum of Maxentius.
 The software provides a VR environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index.
@@
 Inside the immersive environment, the user will start in front of the Maxentius A stairway.
-The user can move around, using the continuous movement metaphor, and can snap-turn or move around freely, depending on the physical space available around them.
+The user can move around, using the continuous movement metaphor, and can snap-turn to rotate.
 Figure 4: The software we provide together with one of our models. The panel, the miniature, and the actual visualization can be seen.
-There is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a coloring akin to the extended matrix by Demetrescu by touching those buttons with their hands.
+There is a panel with buttons to switch between the three reconstruction models and a toggle to turn on or off a coloring akin to the extended matrix by Demetrescu.
+Users can pick up this panel and carry it through the VR scene to change the representation at any position.
 Switching between models is especially relevant in a research context for comparing several different hypotheses directly while still in the iVR environment.
@@
 To the right of the panel is a miniature of the model.
-By holding their virtual hands inside the miniature and performing a grabbing motion, a vibration of the controller indicates interaction, and they can control the light’s angle of incidence on the model.
-There is a small yellow sphere indicating the relative position of the light source at the miniature.
+By holding their virtual hands inside the miniature and performing a grabbing motion, a vibration of the controller indicates interaction, allowing them to control the angle of the light on the model.
+There is a small yellow sphere indicating the relative position of the light source to the miniature.
 This small sphere is intended to serve as a visualization to ease the intuition of where the light comes from.
-This idiom is supposed to help when analyzing how and where sunlight comes from for the mausoleum.
-For our mausoleum scenario, it would make sense to look into relevant ways to interact with light sources.
+This idiom is intended to aid in analyzing the source and distribution of sunlight for the mausoleum.
+For our mausoleum scenario, it would be beneficial to explore ways to effectively interact with light sources.
 An optimized light interaction system could lead to a more fluent experience, which in turn may help experts justify the different shapes of the dome and the light conditions inside.
@@
 For example, users could be enabled to colorize specific areas of the wall to mark the locations of the numerous drawings on the walls, which were added long after the mausoleum was originally built.
-The type of marker, color coding, or maybe even an acoustic signal could then again be investigated for its effectiveness in finding the different spots.
-Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need for additional hardware.
+The type of marker, color coding, or even an acoustic signal could then be investigated for its effectiveness in locating the different spots.
+Lastly, the software can also be used without a VR headset, utilizing very simple keyboard and mouse controls, allowing it to be viewed without the need for additional hardware.
 We provide the software via a Zenodo upload at [35].
@@
 With this work, we present three digital reconstructions of the Maxentius Mausoleum in Rome and utilize them to assess the value of IA for virtual archaeology.
-The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea and are based on archaeological evidence and scholarly interpretations.
-Such scientific evidence provides us with a sound foundation for our discussion in which we want to advocate for the use of immersive technologies more prominently for research purposes.
-While there exist works that point out the usefulness of immersive environments in Virtual Archaeology [29], they tend to focus on dissemination or reconstruction and not on how to improve the perception of a specific aspect, like the comparison of hypotheses.
-Using the three reconstruction models, we created immersive reconstructions in the form of an interactive VR (iVR) environment.
+The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea, based on archaeological evidence and scholarly interpretations.
+Such scientific evidence provides a sound foundation for our discussion, in which we aim to advocate for the more prominent use of immersive technologies to support archaeological research.
+While there exist works that highlight the usefulness of immersive environments in virtual archaeology [29], they tend to focus on dissemination or reconstruction rather than improving the perception of a specific aspect, such as the comparison of hypotheses.
+From a historical and archaeological perspective, the immersive reconstruction of Maxentius’ mausoleum provides important insights into the changing relationship between light and space, a topic much debated in discussions on the development of Late Antique architecture.
+As an outstanding example of a new type of two-storeyed mausolea combining a dark crypt with light-flooded upper floors, it adds a completely new dimension of including light to funeral architecture that was taken up and further developed in the fourth century by church-attached imperial mausolea [Hesberg1992].
+The graded lightning seen here became an important feature of late antique interiors, extending from funeral to palatial [gunter1968wand] and ecclesiastical [deBlaauw2010] architecture.
+An immersive reconstruction of Maxentius’ mausoleum complex will explore this dimension of light and lightning in a systematical manner and, at the same time, offer a promising starting point to take the general discussion of this important feature of late antique architecture to a new level.
+Thus, using the three reconstruction models, we created immersive reconstructions in the form of an interactive VR (iVR) environment.
 This iVR environment serves as proof of concept of the potential of IA in archaeological analysis.
-We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatio-temporal data.
-Our iVR environment allows users to manipulate lighting conditions, toggle between showing a color coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor.
-They use their hands to change the lighting and gain haptic feedback when using buttons inside the virtual space.
+We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation.
+Our environment enables users to manipulate lighting conditions, toggle between color-coded representations of reconstruction certainty and grey, place spatial markers, and navigate with a continuous movement metaphor.
+They use their hands to adjust the lighting and gain haptic feedback when using buttons within the virtual space.
 We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension [11].
-We chose VR because the ability to interact with the reconstruction in an immersive environment supports a deeper understanding of spatial relationships.
-This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy [28].
+We chose VR because it enables interaction with the reconstruction in an immersive environment, supporting a deeper understanding of spatial relationships.
+This finding is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy [28].
 The color coding we use builds upon established practices in virtual archaeology to represent reliability in reconstructions [9].
@@
 Our objective was not to provide comprehensive coverage but to demonstrate a number of possibilities as an example.
-The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching through them on demand.
+The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching between them on demand.
 Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those that are not known but can be derived from similar architecture.
-Through the miniature, users can see how the light source’s position changes relative to the reconstruction.
-The general implications of this work are significant for both archaeological methodology and public engagement.
-It emphasizes a number of ways to interact with and analyze archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from.
-The ability to explore spatial relationships, lighting conditions, and levels of certainty in an immersive environment with embedded analysis information could lead to new research questions.
-Due to the character of this work, there are a number of limitations to consider.
+Through the miniature, users can see how the position of the light source changes relative to the reconstruction.
+Due to the nature of this work, several limitations should be considered.
 As for the models we created, as is usual in the domain, they also have parts that are more interpretative than others.
@@
 Although multi-modality is important, only haptic feedback and a visual representation of the models are available.
-Additionally, the use of VR hardware limits the accessibility for some users.
-While we do offer simple controls that allow the software without VR hardware, the immersion factor that comes with VR is lost.
-Also, the considerations we make are a number of potential intersections of the two fields without taking deep considerations of how the design space of the different methods, like interaction design, would have to be applied.
-Thus, future work should include a variety of aspects.
-For one, a unified framework for creating immersive experiences would pose an asset.
-Such a framework should cover the inclusion and on-demand availability of historical background inside an immersive environment or the sonification of information, both of which have benefits.
-It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time and date.
+Also, the use of VR hardware limits accessibility for some users.
+While we do offer simple controls that allow the software to be used without VR hardware, the immersion factor that comes with VR is lost.
+Additionally, the considerations we make involve several potential intersections of the two fields, without taking into account the in-depth considerations of how the design space of different methods, such as interaction design, would need to be applied.
+For future work, a unified framework for creating immersive experiences would pose an asset.
+Such a framework should encompass the inclusion and on-demand availability of historical context within an immersive environment.
+This could happen in the form of integrated analysis tools, such as classifiers for identifying anachronistic components within a digitization.
+The sonification of information may also be of interest.
+It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time, date, and seasonal characteristics, which was out of the scope of this work.
 However, not only software is required.
 A relevant direction is to assess specific and established workflows from IA and how they can be transferred and used in virtual archaeology.
-It could also be valuable to compare the effectiveness of immersive environments with traditional methods of archaeological data analysis.
+It could also be valuable to compare the effectiveness of immersive environments with traditional methods of archaeological data analysis by evaluating the performance of experts on related tasks, such as identifying archaeological units.
 Lastly, it would be interesting to see the effect of new metaphors for representing inherent information, such as levels of certainty, volumes of objects, sizes, usage, or chronological embedding.
-Therefore, there are a great number of opportunities in the future in this intersection, which should be discovered and evaluated together with domain experts to gain further insights into the usage of the Maxentius Mausoleum specifically and cultural heritage objects in general.
+Therefore, there are numerous opportunities in the future at this intersection, which should be explored and evaluated together with domain experts to gain further insights into the use of the Maxentius Mausoleum, specifically, and cultural heritage objects in general.
 VI Conclusion
-In this work, we present three reconstruction hypotheses of the Maxentius Mausoleum and the literature and typology that inspired them.
-In addition, we have created an immersive environment in which those reconstructions can be viewed and explored.
-The environment comes with a small number of integrated interaction and navigation possibilities to provide not only the models but also the means to explore them.
-Our second contribution is the exploration of the potential for Virtual Archaeology that Immersive Analytics as a research direction yields.
+In this work, we present three reconstruction hypotheses of the Maxentius Mausoleum, along with the literature and typology that inspired them.
+We created 3D models of those reconstruction hypotheses and provide them openly.
+In addition, we have created an immersive environment in which these reconstructions can be viewed and explored through a small number of integrated interaction and navigation options, either as VR or desktop applications.
+Furthermore, we explore the potential for virtual archaeology that immersive analytics as a research direction yields.
 We advocate the shift from ever-better representations, recording, and storage of cultural heritage items to tailored immersive experiences that improve accessibility and insight into existing data.
-The direction is to look at how virtual archaeology studies can be optimized by identifying the specific requirements of a given problem, like comparing different monument reconstruction interpretations and finding out how to support that.
+The goal is to look at how virtual archaeology studies can be optimized by identifying the specific requirements of a given problem, such as comparing different monument reconstruction interpretations and investigating support options.
 For argumentation, we outline the possibilities of engaging representations of reconstructions and interactive representations of hypotheses.
-We exemplify the concept of integrating IA with our current research on the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding its construction and use.
-In our future research, we want to look into how the workflows from IA can be integrated into digital reconstruction.
-We like to combine existing modeling workflows for cultural heritage objects and advancements in data representation in immersive environments.
-We will continue to utilize the Maxentius Mausoleum for that and branch out to other objects, where appropriate.
+We exemplify the concept of integrating IA with our current research on the Maxentius Mausoleum, where these concepts will facilitate the analysis of hypotheses regarding its construction, use, and the relationship between light and space within it.
 Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.